# Pretrained deep models outperform GBDTs in Learning-To-Rank under label scarcity

### v4

@@ Abstract
On tabular data, a significant body of literature has shown that current deep learning (DL) models perform at best similarly to Gradient Boosted Decision Trees (GBDTs), while significantly underperforming them on outlier data Gorishniy et al. (2021); Rubachev et al. (2022); McElfresh et al. (2023). However, these works often study idealized problem settings which may fail to capture complexities of real-world scenarios. We identify a natural tabular data setting where DL models can outperform GBDTs: tabular Learning-to-Rank (LTR) under label scarcity. Tabular LTR applications, including search and recommendation, often have an abundance of unlabeled data, and scarce labeled data. We show that DL rankers can utilize unsupervised pretraining to exploit this unlabeled data. In extensive experiments over both public and proprietary datasets, we show that pretrained DL rankers consistently outperform GBDT rankers on ranking metrics—sometimes by as much as $38\%$—both overall and on outliers.
@@ 1 Introduction
Figure 1: Main result (Simulated crowdsourcing of labels): We compare NDCG ($\uparrow$) as we sweep the percentage of labeled training query groups (QGs); labeled QGs have relevance scores for every element of the QG. For small enough fractions of labeled query groups, pre-trained DL rankers (sometimes significantly) outperform both GBDTs (represents the best of supervised GBDTs and semi-supervised GBDTs) and non-pretrained DL rankers (represents the best of tabular ResNet (Gorishniy et al., 2021), DeepFM (Guo et al., 2017), and DCNv2 (Wang et al., 2021)). To the left of the black dotted line, pretrained rankers perform the best (statistically significant at $p=0.05$ level). Points are averages over 3 trials.
A natural question is whether deep models can similarly use unsupervised pretraining to outperform GBDTs on the tabular LTR problem. To this end, many papers have studied unsupervised pretraining techniques—both using deep learning and GBDTs—for tabular data Yoon et al. (2020); Ucar et al. (2021); Bahri et al. (2021); Verma et al. (2021); Majmundar et al. (2022); Duh & Kirchhoff (2008); Pseudo-Label (2013). However, none of these methods have shown convincingly that deep learning (with or without unsupervised pretraining) consistently outperforms GBDTs (semi-supervised or supervised); at best, deep methods appear to achieve roughly the same performance as GBDTs (Jeffares et al., 2023; Rubachev et al., 2022; Gorishniy et al., 2021; Shwartz-Ziv & Armon, 2022; Grinsztajn et al., 2022). Furthermore, recent work showed that GBDTs are significantly more robust against outliers in the dataset than deep learning models McElfresh et al. (2023). We note that most of these works study idealized problem settings which may fail to capture complexities of real-world scenarios such as label scarcity and noisy data. We hypothesize that pretrained deep models can beat GBDTs in settings that more accurately reflect real-world challenges.
@@
Contributions: (1) We demonstrate for the first time that unsupervised pretraining can produce deep models that outperform GBDTs in ranking. Specifically, pretrained deep rankers can achieve up to 38% higher NDCG (normalized discounted cumulative gain) score (Burges, 2010) than GBDTs when labels are scarce (Figure 1). For instance, Figure 1 shows on three datasets that as the labels get scarcer, GBDTs start to substantially underperform pretrained deep rankers (experimental setup in Section 4.1.2).
@@
(3) We demonstrate that when labels are scarce, pretrained deep rankers can perform significantly better than GBDTs on outlier query groups, boosting NDCG by as much as 80% (Figure 8, Section A.3.4). This is perhaps surprising, given prior results in the label-rich regime showing that GBDTs are significantly more robust to data with irregularities than deep rankers (McElfresh et al., 2023).
@@
Related work: We focus on the traditional LTR setting where the features are all numeric (tabular data). In this setting, gradient boosted decision trees (GBDTs) (Friedman, 2001) are the de-facto models, and deep models have yet to outperform them convincingly (Qin et al., 2021; Joachims, 2006; Ai et al., 2019; Bruch et al., 2019; Ai et al., 2018; Pang et al., 2020; McElfresh et al., 2023). This includes factorization machines (Rendle, 2010), wide & deep architectures (Cheng et al., 2016), and similar architectures designed for recommendation systems (Guo et al., 2017; Wang et al., 2021; Lian et al., 2018; Naumov et al., 2019; Qu et al., 2016). While our paper is architecture agnostic, we perform comparisons of pretrained models against non-pretrained DeepFM and DCNv2 models to show that pretraining is needed to outperform GBDTs consistently. Self-supervised learning (SSL) or unsupervised pretraining has improved overall performance and robustness to noise (Hendrycks et al., 2019) in settings where there is a significant source of unlabeled data like text (Devlin et al., 2018; Howard & Ruder, 2018; Kitaev et al., 2018; Liu et al., 2017; Song & Roth, 2015) and images (Chen et al., 2020; Chen & He, 2021; Grill et al., 2020; Ge et al., 2021; Xie et al., 2020).
@@
Inspired by the success of SSL in images and text, many unsupervised pretraining tasks have been proposed for tabular data. We describe some of the most well-known methods. SubTab (Ucar et al., 2021) proposes to train an autoencoder to reconstruct a table using a subset of the columns. VIME-self (Yoon et al., 2020) is a similar method. SCARF and DACL+ (Bahri et al., 2021; Verma et al., 2021) propose augmentations for the SimCLR pretraining task (Chen et al., 2020) that work for the tabular setting. RegCLR (Wang et al., 2022a) is a method for extracting tables from images, which is not directly applicable to our setting. We comprehensively evaluate/compare SubTab, SCARF, DACL+, VIME-self, SimCLR-Rank, SimSiam (Chen & He, 2021), and self-supervised approaches for GBDTs in Section 3.1 and find that SimCLR-Rank along with SimSiam (Chen & He, 2021) are the best for the tabular LTR setting. We provide a detailed supplemental discussion of related works in the Appendix A.1.
@@ 2 Learning-To-Rank problem and its metrics
In the LTR problem, samples are query groups (QGs) consisting of a query (e.g. shopping query) and $L$ potentially relevant items (e.g. products) for this query. The $i$-th relevant item here is represented by a feature vector $\bm{x}_{i}\in\mathbb{R}^{d}$, which captures information about this query-item pair, subsuming any query related features. In labeled data, such as training or testing, we also have an associated scalar relevance label $y_{i}$, which could be a binary, ordinal, or real-valued measurement of the relevance of this item for this query (Qin et al., 2021). The objective is to learn a function that ranks these $L$ items such that the items with highest true relevance are ranked at the top.
Most LTR algorithms formulate the problem as learning a scoring function $f_{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ that maps the feature vector associated with each item to a score, and then ranking the items by sorting the scores in descending order. To measure the quality of a ranking induced by our scoring function $f_{\theta}$, a commonly-used (per-QG) metric is NDCG ($\uparrow$):
@@
where $\pi_{f_{\theta}}:[L]\to[L]$ ($[L]\triangleq\{1,\dots,L\}$) is a ranking of the $L$ elements induced by the scoring function $f_{\theta}$ on $\{\bm{x}_{i}\}_{i=1}^{L_{k}}$ (i.e. $\pi(i)$ is the rank of the $i$-th item), while $\pi^{*}$ is the ideal ranking induced by the true relevance labels $\{y_{i}\}_{i=1}^{L}$, and discounted cumulative gain (DCG) is defined as $\text{DCG}(\pi,\{y_{i}\}_{i=1}^{L})=\sum_{i=1}^{L}\frac{2^{y_{i}}-1}{\text{log}_{2}(1+\pi(i))}.$ Typically, a truncated version of NDCG is used that only considers the top-$k$ ranked items, denoted as NDCG@$k$. In the rest of our paper, we will refer to NDCG@5 as NDCG; this will be our evaluation metric.
@@ 2.1 Outlier-NDCG for outlier performance evaluation
Motivated by a recent observation that GBDTs outperform DL models in datasets with irregularities McElfresh et al. (2023), we also study their performance on outliers. In interactive ML systems like search, performing well on outlier queries is particularly valuable as it empowers users to search for more outlier queries, which in turn allows the modeler to collect more data and improve the model. To this end, we evaluate the outlier performance by computing the average NDCG on outlier queries groups and refer to it as Outlier-NDCG for brevity. In practice, some outlier queries may already be known, and the modeler can define the outlier datasets accordingly, like in our proprietary dataset. For example, since industry data pipelines often have missing data/features one could identify samples with missing features as outliers. When outliers are unknown, detecting them (especially in higher-dimensional data) is challenging. Hence, we create a heuristic algorithm by assuming outliers are rare values that are separated from the bulk of the data. Using it we systematically selected outlier query groups for the public datasets that we use. More details about the outliers are provided in Section A.2.2. Separately measuring the performance on the outliers is justified by our later observation that performance on the full dataset and the outliers can significantly differ (Tables 2 and 3).
@@ 3 Design: Unsupervised pretraining for LTR
In this section, we systematically evaluate different strategies for utilizing unlabeled data in pretraining of tabular LTR models, including a novel LTR-specific pretraining loss, SimCLR-Rank. We ultimately find that the top pretraining strategies are SimCLR-Rank or SimSiam, but the best choice among these depends on the dataset.
@@
Semi-supervised learning for GBDTs: While to the best of our knowledge there are no methods to perform unsupervised pretraining in GBDTs, one can use semi-supervised methods like consistency regularization (Jeong & Shin, 2020; Sajjadi et al., 2016; Miyato et al., 2018; Oliver et al., 2018; Berthelot et al., 2019), pseudo-labeling/self-training (Rizve et al., 2021; Pseudo-Label, 2013; Shi et al., 2018; Yarowsky, 1995; McClosky et al., 2006), and PCA (Duh & Kirchhoff, 2008) to address limited size of labeled data. In consistency regularization, one increases the size of the training set by using data augmentations. In pseudo-labeling/self-training, a model trained on the available labeled data is used to label unlabeled data. Then, the final model is trained on the resulting “labeled” dataset. In PCA, one projects the features along directions of maximum variability in the union of labeled and unlabeled datasets. Since prior work has found that consistency regularization decreases the performance of GBDTs (Rubachev et al., 2022), we evaluate only pseudo-labeling and PCA.
@@
Pretraining methods for tabular data: There have been many unsupervised pretraining methods proposed for tabular data, which can be applied to the tabular LTR setting. In this paper, we will evaluate (1) SCARF (Bahri et al., 2021) and (2) DACL+ (Verma et al., 2021) which are contrastive learning methods (Chen et al., 2020), (3) VIME-self (Yoon et al., 2020) and (4) SubTab (Ucar et al., 2021) which are autoencoder based methods. These four methods are the most commonly-evaluated baselines in the pretraining literature for tabular data (Ucar et al., 2021; Hajiramezanali et al., 2022; Levin et al., 2022; Majmundar et al., 2022).
@@
Domain-agnostic pretraining methods: In the image setting, SimCLR (Chen et al., 2020) and SimSiam (Chen & He, 2021) are two widely-adopted unsupervised pretraining methods. They rely on the idea that two slightly augmented variations of a data point should have similar representations. This simple idea generalizes across domains, so we apply it to tabular LTR. We discuss them below:
@@
SimCLR (Chen et al., 2020): In a batch of $B$ query-groups, the feature vector $\bm{x}_{q,i}$ of $q$-th query group and its $i$-th item is stochastically augmented twice to obtain $\bm{x}_{q,i}^{(0)}$ and $\bm{x}_{q,i}^{(1)}$, which are called a positive pair. In this paper, we use combinations of the following two augmentation strategies for all domain-agnostic pretraining methods: (a) Zeroing: Zero out features selected randomly and i.i.d. with fixed probability. (b) Gaussian noise: Add Gaussian noise of a fixed scale to all features. Further details are given in Section A.2.3.
@@
Figure 2: In SimCLR and its variants, positive pairs, or augmented versions of the same data sample, are trained to have similar embeddings. Positive-negative pairs, or augmented samples originating from two different data points or classes, are trained to have distant embeddings. In vanilla SimCLR, each positive pair is contrasted with all other items in the batch, denoted by the data points contained in the brown dashed line above. In SimCLR-Rank, each positive pair is contrasted with only the items in the same Query Group (QG), denoted by the data points inside the red solid line above. By limiting the negatives to a smaller set of harder-to-contrast relevant items, SimCLR-Rank is faster and scales to larger datasets.
Pretraining method for LTR: Motivated by (a) SimCLR’s high complexity and (b) the efficacy of contrasting with hard negatives, we propose SimCLR-Rank, an LTR-specific alternative. Recall that the SimCLR loss for an item in equation 2 uses all the other items in the batch as negatives. Good pretraining strategies usually exploit the structure of the data, e.g. text sequences motivates masked token prediction Devlin et al. (2018). Studying the query group structure of LTR data, we notice that high-quality hard negatives are freely available as the other items from the same query group. These items are retrieved and deemed potentially similarly relevant for this query by an upstream retrieval model, making these items harder negatives than other items in the batch. So, we propose SimCLR-Rank, which modifies SimCLR to contrast only with the other items in the same query group, as illustrated in Figure 2. Formally, we modify the per-augmentation loss $\ell^{(a)}_{q,i}$ as:
@@
The SimCLR-Rank loss has a time/space complexity of $O(BL^{2})$, making it faster than SimCLR’s $O(B^{2}L^{2})$ complexity, and enjoying SimSiam’s linear complexity in $B$. Next, we empirically study these methods.

### v5

@@ Abstract
On tabular data, a significant body of literature has shown that current deep learning (DL) models perform at best similarly to Gradient Boosted Decision Trees (GBDTs), while significantly underperforming them on outlier data Gorishniy et al. (2021); Rubachev et al. (2022); McElfresh et al. (2023). However, these works often study problem settings which may not fully capture the complexities of real-world scenarios. We identify a natural tabular data setting where DL models can outperform GBDTs: tabular Learning-to-Rank (LTR) under label scarcity. Tabular LTR applications, including search and recommendation, often have an abundance of unlabeled data, and scarce labeled data. We show that DL rankers can utilize unsupervised pretraining to exploit this unlabeled data. In extensive experiments over both public and proprietary datasets, we show that pretrained DL rankers consistently outperform GBDT rankers on ranking metrics—sometimes by as much as $38\%$—both overall and on outliers.
@@ 1 Introduction
Figure 1: Main result (Simulated crowdsourcing of labels): We compare NDCG ($\uparrow$) as we sweep the percentage of labeled training query groups (QGs); labeled QGs have relevance scores for every element of the QG. For small enough fractions of labeled query groups, pre-trained DL rankers (sometimes significantly) outperform both GBDTs (representing the best of supervised GBDTs and semi-supervised GBDTs) and non-pretrained DL rankers (representing the best of tabular ResNet (Gorishniy et al., 2021), DeepFM (Guo et al., 2017), and DCNv2 (Wang et al., 2021)). To the left of the black dotted line, pretrained rankers perform the best (statistically significant at $p=0.05$ level). Points are averages over 3 trials.
A natural question is whether deep models can similarly use unsupervised pretraining to outperform GBDTs on the tabular LTR problem. To this end, many papers have studied unsupervised pretraining techniques—both using deep learning and GBDTs—for tabular data Yoon et al. (2020); Ucar et al. (2021); Bahri et al. (2021); Verma et al. (2021); Majmundar et al. (2022); Duh & Kirchhoff (2008); Pseudo-Label (2013). However, none of these methods have shown convincingly that deep learning (with or without unsupervised pretraining) consistently outperforms GBDTs (semi-supervised or supervised); at best, deep methods appear to achieve roughly the same performance as GBDTs (Jeffares et al., 2023; Rubachev et al., 2022; Gorishniy et al., 2021; Shwartz-Ziv & Armon, 2022; Grinsztajn et al., 2022). Furthermore, recent work showed that GBDTs are significantly more robust against outliers in the dataset than deep learning models McElfresh et al. (2023). We note that most of these works study idealized problem settings which may fail to capture complexities of real-world scenarios such as label scarcity and noisy data. Prior work in other domains show that pretrained deep models can outperform purely supervised models Chen et al. (2020) and former models are more robust to noise Hendrycks et al. (2019). Therefore, we hypothesize that pretrained deep models can outperform GBDTs in tabular settings that more accurately reflect real-world challenges.
@@
Contributions: (1) We demonstrate for the first time that unsupervised pretraining can produce deep models that outperform GBDTs in ranking. Specifically, pretrained deep rankers can achieve up to 38% higher NDCG (normalized discounted cumulative gain) score (Burges, 2010) than GBDTs when labels are scarce (Figure˜1). For instance, Figure˜1 shows on three datasets that as the labels get scarcer, GBDTs start to substantially underperform pretrained deep rankers (experimental setup in Section˜4.1.2).
@@
(3) We demonstrate that when labels are scarce, pretrained deep rankers can perform significantly better than GBDTs on outlier query groups, boosting NDCG by as much as 80% (Figure˜8, Section˜A.3.4). This is perhaps surprising, given prior results in the label-rich regime showing that GBDTs are significantly more robust to data with irregularities than deep rankers (McElfresh et al., 2023).
@@
Related work: We focus on the traditional LTR setting where the features are all numeric (tabular data). In this setting, gradient boosted decision trees (GBDTs) (Friedman, 2001) have historically been the de-facto models, though there is currently great interest in applying deep neural networks to tabular data. Borisov et al. (2022) categorize existing techniques for using deep neural networks over tabular data into four types: (1) Encoding-based methods such as VIME (Yoon et al., 2020), SCARF (Bahri et al., 2021), IGTD (Zhu et al., 2021), and SuperTML (Sun et al., 2019); (2) Novel hybrid architectures such as DeepFM (Guo et al., 2017), xDeepFM (Lian et al., 2018), and many others Cheng et al. (2016); Frosst & Hinton (2017); Ke et al. (2018; 2019); Popov et al. (2019); Luo et al. (2020); Liu et al. (2020); Ivanov & Prokhorenkova (2021); Luo et al. (2021); (3) Transformer-based architectures including SAINT (Somepalli et al., 2021), TabNet (Arik & Pfister, 2021), TabTransformer (Huang et al., 2020), and ARM-Net Cai et al. (2021)); (4) Regularized DNNs (Shavitt & Segal, 2018; Kadra et al., 2021).
@@
Within these categories, proposed architectures include factorization machines (Rendle, 2010), wide & deep architectures (Cheng et al., 2016), and similar architectures designed for recommendation systems (Guo et al., 2017; Wang et al., 2021; Lian et al., 2018; Naumov et al., 2019; Qu et al., 2016). While our paper is architecture agnostic, we perform comparisons of pretrained models against non-pretrained DeepFM Guo et al. (2017) and DCNv2 Wang et al. (2021) models to show that pretraining is needed to outperform GBDTs consistently. Self-supervised learning (SSL) or unsupervised pretraining has improved overall performance and robustness to noise (Hendrycks et al., 2019) in settings where there is a significant source of unlabeled data like text (Devlin et al., 2018; Howard & Ruder, 2018; Kitaev et al., 2018; Liu et al., 2017; Song & Roth, 2015) and images (Chen et al., 2020; Chen & He, 2021; Grill et al., 2020; Ge et al., 2021; Xie et al., 2020).
@@
Inspired by the success of SSL in images and text, many unsupervised pretraining tasks have been proposed for tabular data. We describe some of the most well-known methods. SubTab (Ucar et al., 2021) proposes to train an autoencoder to reconstruct a table using a subset of the columns. VIME-self (Yoon et al., 2020) is a similar method. SCARF and DACL+ (Bahri et al., 2021; Verma et al., 2021) propose augmentations for the SimCLR pretraining task (Chen et al., 2020) that work for the tabular setting. RegCLR (Wang et al., 2022a) is a method for extracting tables from images, which is not directly applicable to our setting.
A smaller class of approaches has studied regularized DNNs, finding that they require limited changes to perform well on tabular data Shavitt & Segal (2018); Kadra et al. (2021). However, the most competitive of these methods require extensive hyperparameter optimization over a 14-dimensional grid Kadra et al. (2021).
While the body of work studying neural network architectures for tabular data is extensive, deep models have yet to outperform GBDTs convincingly (Qin et al., 2021; Joachims, 2006; Ai et al., 2019; Bruch et al., 2019; Ai et al., 2018; Pang et al., 2020; McElfresh et al., 2023). In this work, we consider encoding-based methods, which alter the representations over which a deep model can learn to rank; we do not focus on introducing new algorithms in this work, but rather evaluate mostly existing methods applied to tabular data. We comprehensively evaluate seven representative DL methods: SubTab, SCARF, DACL+, VIME-self, SimCLR-Rank, SAINT, SimSiam (Chen & He, 2021), and self-supervised approaches for GBDTs in Section˜3.1. Our results show that multiple pretrained deep rankers can consistently outperform GBDTs in the label-scarce tabular LTR setting, with SimCLR-Rank and SimSiam (Chen & He, 2021) generally performing the best. We provide a detailed supplemental discussion of related works in the Appendix A.1.
@@ 2 Learning-To-Rank problem and its metrics
In the LTR problem, samples are query groups (QGs) consisting of a query (e.g. shopping query) and $L$ potentially relevant items (e.g. products) for this query. Although an LTR dataset contains multiple QGs, for the sake of simplicity, the notation in this section only handles a single QG. The $i$-th relevant item in this QG is represented by a feature vector $\bm{x}_{i}\in\mathbb{R}^{d}$, which captures information about this query-item pair, subsuming any query related features. If and only if the QG is labeled, such as the ones used in supervised training or testing, we also have an associated scalar relevance label $y_{i}$, which could be a binary, ordinal, or real-valued measurement of the relevance of this item for this query (Qin et al., 2021). The objective is to learn a function that ranks these $L$ items such that the items with highest true relevance are ranked at the top.
Most LTR algorithms formulate the problem as learning a scoring function $f_{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ that maps the feature vector associated with each item to a score, and then ranking the items by sorting the scores in descending order. To measure the quality of a ranking induced by our scoring function $f_{\theta}$ on a labeled test QG, a commonly-used (per-QG) metric is NDCG (normalized cumulative discounted gain) ($\uparrow$):
@@
where $\pi_{f_{\theta}}:[L]\to[L]$ ($L]\triangleq\{1,\dots,L\}$) is a ranking of the $L$ elements induced by the scoring function $f_{\theta}$ on $\{\bm{x}_{i}\}_{i=1}^{L}$ (i.e. $\pi(i)$ is the rank of the $i$-th item), while $\pi^{*}$ is the ideal ranking induced by the true relevance labels $\{y_{i}\}_{i=1}^{L}$, and discounted cumulative gain (DCG) is defined as $\text{DCG}(\pi,\{y_{i}\}_{i=1}^{L})=\sum_{i=1}^{L}\frac{2^{y_{i}}-1}{\text{log}_{2}(1+\pi(i))}.$ Here, ($\uparrow$) indicates that larger metric values are preferred. Note that for summarizing this metric over the full test set, we simply take the average across all the QGs in the set. Typically, a truncated version of NDCG is used that only considers the top-$k$ ranked items, denoted as NDCG@$k$. In the rest of our paper, we will refer to NDCG@5 as NDCG; this will be our evaluation metric.
@@ 2.1 Outlier-NDCG for outlier performance evaluation
Motivated by a recent observation that GBDTs outperform DL models in datasets with irregularities McElfresh et al. (2023), we also study their performance on outliers. In interactive ML systems like search, performing well on outlier queries is particularly valuable as it empowers users to search for more outlier queries, which in turn allows the modeler to collect more data and improve the model. To this end, we evaluate the outlier performance by computing the average NDCG on outlier queries groups and refer to it as Outlier-NDCG for brevity. In practice, some outlier queries may already be known, and the modeler can define the outlier datasets accordingly, like in our proprietary dataset. For example, since industry data pipelines often have missing data/features one could identify samples with missing features as outliers. When outliers are unknown, detecting them (especially in higher-dimensional data) is challenging. Hence, we create a heuristic algorithm by assuming outliers are rare values that are separated from the bulk of the data. Using it we systematically selected outlier query groups for the public datasets that we use. More details about the outliers are provided in Section˜A.2.2. Separately measuring the performance on the outliers is justified by our later observation that performance on the full dataset and the outliers can significantly differ (Tables˜2 and 3).
@@ 3 Design: Unsupervised pretraining for LTR
Qualitatively, the ranking relevance of tabular records to a query is typically determined by a combination of fields. For example, if a person searches for a product (say a pair of sandals), we can think of the fields in the LTR problem as representing properties of the results, such as material, color, and price. Qualitatively, we expect that if a pair of records are the same except for small differences between fields (e.g., missing a field), then the two products should generally have similar rankings in the LTR problem. In this section, we discuss how existing methods fail to capture this qualitative intuition.
@@
Semi-supervised learning for GBDTs: While to the best of our knowledge there are no methods to perform unsupervised pretraining in GBDTs, one can use semi-supervised methods like consistency regularization (Jeong & Shin, 2020; Sajjadi et al., 2016; Miyato et al., 2018; Oliver et al., 2018; Berthelot et al., 2019), pseudo-labeling/self-training (Rizve et al., 2021; Pseudo-Label, 2013; Shi et al., 2018; Yarowsky, 1995; McClosky et al., 2006), and PCA (Duh & Kirchhoff, 2008) to address limited size of labeled data. In consistency regularization, one increases the size of the training set by using data augmentations. In pseudo-labeling/self-training, a model trained on the available labeled data is used to label unlabeled data. Then, the final model is trained on the resulting “labeled” dataset. In PCA, one projects the features along directions of maximum variability in the union of labeled and unlabeled datasets. Since prior work has found that consistency regularization decreases the performance of GBDTs (Rubachev et al., 2022), we evaluate only pseudo-labeling and PCA. None of these methods explicitly ensures that semantically similar records have similar representations in the learned models.
@@
Pretraining methods for tabular data: To address this gap, there have been many unsupervised pretraining methods proposed for tabular data, which can be applied to the tabular LTR setting. In this paper, we will evaluate (1) SCARF (Bahri et al., 2021) and (2) DACL+ (Verma et al., 2021) which are contrastive learning methods (Chen et al., 2020), (3) VIME-self (Yoon et al., 2020) and (4) SubTab (Ucar et al., 2021) which are autoencoder based methods. The four most commonly-evaluated baselines in the pretraining literature for tabular data are Ucar et al. (2021); Hajiramezanali et al. (2022); Levin et al. (2022); Majmundar et al. (2022). Our results suggest that many of these methods outperform GBDTs in the label-scarce regime (Section˜3.1), in accordance with the main theme of this paper. However, the consistently best-performing methods in our experiments were more basic, domain-agnostic pretraining methods, which can be viewed as more primitive building blocks of some of the above methods. Surprisingly, they use less domain knowledge of tabular data than these more sophisticated methods.
@@
Specifically, in the image setting, SimCLR (Chen et al., 2020) and SimSiam (Chen & He, 2021) are two widely-adopted unsupervised pretraining methods. They use the idea that two variations of the same data point should have similar representations. This idea generalizes across domains, including tabular LTR.
@@
SimCLR (Chen et al., 2020): In a batch of $B$ query-groups, the feature vector $\bm{x}_{q,i}$ of $q$-th query group and its $i$-th item is stochastically augmented twice to obtain $\bm{x}_{q,i}^{(0)}$ and $\bm{x}_{q,i}^{(1)}$, which are called a positive pair. In this paper, we use combinations of the following two augmentation strategies for all domain-agnostic pretraining methods: (a) Zeroing: Zero out features selected randomly and i.i.d. with fixed probability. (b) Gaussian noise: Add Gaussian noise of a fixed scale to all features. Further details are given in Section˜A.2.3.
@@
Figure 2: In SimCLR and its variants, positive pairs, or augmented versions of the same data sample, are trained to have similar embeddings. Positive-negative pairs, or augmented samples originating from two different data points or classes, are trained to have distant embeddings. In vanilla SimCLR, each positive pair is contrasted with all other items in the batch, denoted by the data points contained in the brown dashed line above. In SimCLR-Rank, each positive pair is contrasted with only the items in the same Query Group (QG), denoted by the data points inside the red solid line above.
Pretraining method for LTR: Motivated by (a) SimCLR’s high complexity and (b) the efficacy of contrasting with hard negatives, we additionally propose SimCLR-Rank, an LTR-specific alternative. Recall that the SimCLR loss for an item in equation 2 uses all the other items in the batch as negatives. Good pretraining strategies usually exploit the structure of the data, e.g. text sequences motivates masked token prediction Devlin et al. (2018). Studying the query group structure of LTR data, we notice that high-quality hard negatives are freely available as the other items from the same query group. These items are retrieved and deemed potentially similarly relevant for this query by an upstream retrieval model, making these items harder negatives than other items in the batch. So, we propose SimCLR-Rank, which modifies SimCLR to contrast only with the other items in the same query group, as illustrated in Figure˜2. Formally, we modify the per-augmentation loss $\ell^{(a)}_{q,i}$ as:
@@
Empirically, SimCLR-Rank better separates embeddings within query groups relative to sampling negatives uniformly from each batch (Section˜A.5). The SimCLR-Rank loss has a time/space complexity of $O(BL^{2})$, making it faster than SimCLR’s $O(B^{2}L^{2})$ complexity, and enjoying SimSiam’s linear complexity in $B$. Next, we empirically study these methods.